Medical research has concluded that people experience a common cold roughly two times per year.Assume that the time between colds is normally distributed with a mean of 160 days and a standard deviation of 40 days. a-what is the probability of going 200 or more days between colds ? b- what is the probabilty of going 365 or more days ? c- what is the probability of going between 250 and 320 days? d- what is the probability of getting a cold within 80 days of a previous cold?
Answer by Boreal(15235)

z=(x-mean)/sd
> = (200-160)/40 which is > =+1 or 0.1587.
===================
>=365 days is (365-160)/40=205/40=5.125 or <0.0001 (1.5 x 10^-7)
==================
between 250 and 320 days is z between 2.25 and 4 or 0.0122
<=(80-160)/40 or < = -2 or 0.0228